Hello my name is Aaliyah and I really need help on this problem.
2x-5y=-158
x-y=-40
But it will not show me how to do it or anything and really need it, because i am a 7th grader and i am really smart but just do not get this problem at all, and my math class is all 9th graders, only me and three other 7th grade girls are in this honorable class. So please help. Thankyou very much!!!!!!!!!!!!!!!

Found 3 solutions by nerdybill, feliz1965, stanbon:
Answer by nerdybill(7384)

Given:
2x-5y = -158
x- y = -40
.
Using the "addition method", we start by multiplying the second equation by -5:
2x-5y = -158
-5(x- y) = -5(-40)
.
Giving us:
2x-5y = -158
-5x+5y = 200
.
Now we ADD the two equation together:
2x-5y = -158
-5x+5y = 200
----------------
-3x = 42
.
Finally, we divide both sides by -3:
-3x = 42
x = -14
.
To find y, we plug the above into either equation -- I'll use:
x- y = -40
-14- y = -40
-y = -26
y = 26
.
Solution: (x,y) = (-14, 26)


Answer by feliz1965(151)

2x-5y=-158
x-y=-40
Your class has been introduced to a system of two equations in two variables.
I will use the substitution method to solve for x and y. It is one of 4 ways to solve this problem.
You can also solve this question using the following methods:
1-addition method
2-graphing method
3-matrix algebra

2x-5y=-158
x-y=-40

I will isolate x in the equation x - y = -40 and then replace x in the first equation to find the value of y.
x - y = -40
x = y - 40
Now, I will go back to the first equation and plug that value of x.
2(y - 40)-5y = -158
2y - 80 - 5y = -158
-3y - 80 = -158
-3y = 80 - 158
-3y = -78
y = -78/-3
y = 26
I just found the value of y.
Did you see what I did to find y?
We now need to find x.
I will plug that value of y I just found into EITHER of the original equations given and solve for x.
I selected the equation x - y = -40 but I could have chosen the other equation. It is our choice.
We know that y = 26, right?
So, x - 26 = -40
x = 26 - 40
x = -14
The solution to this system of equations is the point (-14, 26). It is at this point where the two equations cross or meet on the coordinate plane.
The value of x = -14 and y = 26.
Not bad for a sub teacher, right?




Answer by stanbon(75887)

There are several ways to solve the system: elimination, graphing, and substitution are three of the ways.
-------------------
I'll use elimination:
2x-5y=-158
x-y=-40
-----------------
Multiply thru the 2nd equation by 2 to set it up for elimination of the x terms.
2x - 5y = -158
2x - 2y = -80
---------------------
Subtract the top equation from the bottom equation to solve for "y":
3y = 78
y = 26
---------------
Substitute that value of "y" into x - y = -40 to solve for "x":
x - 26 = -40
x = - 14
------------------
The solution is (-14,26)
================================
Cheers,
Stan H.